OpenStudy (anonymous):

What is the significance of the points found by equating first derivative to the second derivative

OpenStudy (amistre64):

second derivative helps to weed out the points of inflection

OpenStudy (amistre64):

it also helps to define the concavity of the points

OpenStudy (anonymous):

Can you give me an article on a site to study on these points ?

OpenStudy (amistre64):

nothing comes to mind. The webpages presented at the bottom of the page are good resources

OpenStudy (anonymous):

Ok so the points on inflection of x^3 - x would be 3x^2 - 1 = 6x right ?

OpenStudy (amistre64):

is f(x) = x^3 -x ?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

the f'(x) = 3x^2-1 3x^2-1 =0 will give us critical points to check; and there doesnt seem to be any value to make it go undetermined ...

OpenStudy (amistre64):

3x^2-1 = 0 3x^2 = 1 x^2 = 1/3 x = +- sqrt(1/3) is a point to notice right?

OpenStudy (anonymous):

yes i know that but what about the points where the graph changes from convexity to concavitivty..? We get it by equation 1st and 2nd derivative like this right ? 3x^2 - 1 = 6x

OpenStudy (amistre64):

no; we get it by finding critical points of f' and f''

OpenStudy (amistre64):

f'(x) = 3x^2 -1 f''(x) = 6x at 6x=0 we have inflection

OpenStudy (anonymous):

Ok

OpenStudy (amistre64):

f' gives us points that may or maynot be inflections; since an inflection can hide behind an f'=0 but, if f''=0 we have a higher degree of certainty that its an inflection; but the best option is to test it out by noticing if the sign stays the same or changes from left to right

OpenStudy (amistre64):

6x = 0 when x = 0 <.........0..........> - + sign changes which means concavity changes; it has to be an inflection